COMPUTATIONAL INVESTIGATION OF THE T_{1}(n, π^{∗}) STATE OF 2-CYCLOHEXEN-1-ONE

\underline{MICHAEL O. McANALLY} and STEPHEN DRUCKER, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702.

We have used several computational methods, including TDDFT and EOM–EE–CCSD, to determine the equilibrium geometry and harmonic vibrational frequencies of 2-cyclohexen-1-one (CHO) in its T_{1}(n, π^{∗}) excited state. Atom displacement vectors for the normal modes indicate that the lowest-frequency vibration, ν_{39}, is best described as a ring-twisting motion, whereas ν_{38} is a ring-bending vibration consisting mainly of C-5 displacement toward and away from the plane in which the other heavy atoms lie. These updated descriptions are transposed with respect to those in the previous literature.^{a} The table below shows that the EOM–EE–CCSD harmonic frequencies generally agree well with fundamentals obtained spectroscopically.^{b} In particular, the ν_{39} frequency determined by EOM–EE–CCSD is more accurate than the TDB3LYP result, with errors of +2% and +20%, respectively. This outcome is traceable in part to a larger ν_{39} reduced mass calculated by EOM–EE–CCSD, stemming from a less planar O=C–C=C dihedral angle (170.4^{∘} via EOM–EE–CCSD vs. 177.4^{∘} via TDB3LYP).

Low-frequency fundamentals (cm^{−1}) for CHO in its T_{1}(n, π^{∗}) state

Mode  Description  TDB3LYP/cc-pVTZ EOM–EE–CCSD/cc-pVDZ Experiment^{b}
39  ring twist  119.6  101.4  99.5
38  bend (inversion of C-5)  255.9  272.0  253.2
37  C=C twist  306.9  223.0  247.8
36  C=O wag  359.0  340.2  345.5

The success of EOM–EE–CCSD in this application could be due to its ability to describe multiconfigurational wavefunctions within a single-reference formalism.

aT. L. Smithson and H. Wieser, J. Chem. Phys. 73, 2518 (1980); M. Z. M. Rishard, E. A. Brown, L. K. Ausman, S. Drucker, J. Choo, and J. Laane, J. Phys. Chem. A 112, 38 (2008).

bPrevious talk.